OpenStudy (anonymous):

If you are given a formula for f, how do you nd a formula for f(inverse 1)?

OpenStudy (anonymous):

i m frm rajasthan

OpenStudy (anonymous):

take the example f (x) = x + 7 to find the inverse, take f(x) as x and replace x as y in the original equation/function so you get x = y + 7 make y the subject of the formula y = x - 7 That is the inverse

OpenStudy (anonymous):

oh ok